OpenStudy (anonymous):

logx to the base a * log y to the base b= logx to the base b * log y to the base a.please help

hartnn (hartnn):

use this : \(\Large \log_xy = \dfrac{\log y}{\log x}\)

OpenStudy (mathmath333):

\(\Large\rm property~~~~\log_xy = \dfrac{\log_m y}{\log _m x} \) \(\Large\rm \log_a x \times \log_b y \) \(\Large\rm =\dfrac{\log_b x}{\log_b a} \times \dfrac{\log_a y}{\log_a b} \) \(\Large\rm =\dfrac{\log_b x\times \log_a y }{\log_b a \times \log_a b } \) \(\Large\rm =\log_b x \times \log_a y \) \(\Large\rm as~~ \log_b a \times \log_a b= \dfrac{\log_a a}{\log_a b } \times \log_a b =\log_a a=1\)